Two sides and an angle are given below. Determine whether the given information results in one​ triangle, two​ triangles, or no triangle at all. Solve any​ triangle(s) that results
b=8 c=10 B=20 degrees

Answer by Alan3354(69443)

Two sides and an angle are given below. Determine whether the given information results in one​ triangle, two​ triangles, or no triangle at all. Solve any​ triangle(s) that results
b=8 c=10 B=20 degrees
------------
Assuming angle B is opposite side b, etc.:
b/sin(B) = c/sin(C)
8/sin(20) = 10/sin(C)
sin(C) = 10*sin(20)/8
sin(C) =~ 0.4275252
==============
It's a triangle.
Can you do the rest?